How do you solve for a when (a-3)squared = 5-10a?
Answer by checkley77(12844)

(a-3)^2=5-10a
a^2-6a+9=5-10a
a^2-6a+10a+9-5=0
a^2+4a+4=0
(a+2)^2=0
a+=0
a=-2 ans.